OpenStudy (anonymous):

the equation -15x+4y=0 relates the length of a videotape in inches x to its approximate playing time in seconds y. will give medal!!!

OpenStudy (anonymous):

What is your question? You've given a bunch of information but haven't actually asked anything.

OpenStudy (anonymous):

tell whether each relationship is a direct variation and why

OpenStudy (anonymous):

An equation represents direct variation if it is of the form: \[y=kx\] (Which is a special case of y=mx+b with b=0) To find out if our equation is direct variation we need to solve for y: \[-15x+4y=0\]\[4y=15x\]\[y=\frac{ 15 }{ 4 }x\] We can see that his matches the pattern of \[y=kx\] with k=15/4 So, yes, this equation is direct variation. The reason is because the ration between x and y is always the same (in this case 15/4). If we double x we double y. If we triple x we triple y. etc. etc.